OpenStudy (calculusxy):

In a race, Alice and Brenda started at the same time and ran with constant spends of 12 km and 20 km per hour, respectively. If Alice crossed the finish line 1 minute after Brenda, how many kilometers long was the race?

OpenStudy (calculusxy):

@jim_thompson5910

OpenStudy (calculusxy):

Since it said "minute" somewhere in the problem, I thought of converting the 12km/hr and 20km/hr into minutes. But I am not sure if that's how I am supposed to start it or if that's even correct.

OpenStudy (calculusxy):

\[\frac{ 12 }{ 60 } = \frac{ 1km }{ 5minutes }\] \[\frac{ 20 }{ 60 } = \frac{ 1 km }{ 3 minute }\]

OpenStudy (ybarrap):

A minute is 1/60 hr, right? So $$ (t+1/60)12=20t $$ Solve for t then plug back in to find the total distance t is in hours Does this make sense

OpenStudy (calculusxy):

So I I just use your formula then I will get it correct?

OpenStudy (calculusxy):

12t + 1/5 = 20t 12t - 12t + 1/5 = 20t - 12t 1/5 = 8t 1/5 / 8 = 8t / 8 1/40 = t

jimthompson5910 (jim_thompson5910):

looks good, now plug that into either side of the original equation to find the distance 20t is the easier side

OpenStudy (calculusxy):

20/40 = 1/2 Okay thanks :)

OpenStudy (calculusxy):

But how did @ybarrap make that equation?

jimthompson5910 (jim_thompson5910):

he used d = r*t d = distance r = rate t = time

OpenStudy (calculusxy):

Also, we are learning about the method called "Backsolving" so if you can help me with that on this problem.

jimthompson5910 (jim_thompson5910):

on the left side r = 12 t is t+1/60 on the right side r = 20 t is just itself: t

jimthompson5910 (jim_thompson5910):